# ORION: Option-Regularized Deep Reinforcement Learning for Cooperative Multi-Agent Online Navigation

### v2

@@ III-D1 Graph-based Encoders
III-D2 Option Decoder $\pi_{\vartheta}$
During multi-agent navigation, an agent may either navigate toward its designated target or assist others by collecting more information that benefits the team, each demanding distinct and sometimes conflicting reasoning. In such cases, adaptive switching between self-directed and cooperative behaviors becomes crucial. To balance individual efficiency and collective performance for improved overall task success, we introduce the option-critic structure [1], where an option denotes a high-level behavioral mode characterized by a termination head function and an intra-option policy, while the critic provides a value-based estimation guiding high-level option/low-level action optimization. Specifically, we augment the policy network with an option decoder, which determines whether to continue the current high-level behavior or terminate and switch to another.
Termination Head $\beta_{\vartheta}$
Given the current state representation feature $\mathbf{s}_{t}$ and the active option $z_{t-1}$ at the last step, we define a termination head function $\beta_{\vartheta}(\mathbf{s}_{t},z_{t-1})$ that estimates the likelihood of terminating the current option at step $t$. Specifically, we first obtain a joint feature representation by combining the current node feature and the embedding of the previous option $h_{t}=\mathbf{s}_{t}+e(z_{t-1}),$ where $e(z_{t-1})\in\mathbb{R}^{d}$ denotes the learnable option embedding. We parameterize the termination function as $\beta_{\vartheta}(\mathbf{s}_{t},z_{t-1})=\sigma\!\big(\mathrm{MLP}_{\vartheta}(h_{t})\big),$ where $\mathrm{MLP}_{\vartheta}(\cdot)$ is a two-layer feed-forward network with ReLU activation, and $\sigma(\cdot)$ denotes the logistic sigmoid. Finally, a Bernoulli sampling $\mathbf{1}_{\text{term}}/\mathbf{1}_{\text{cont}}\sim\mathrm{Bern}(\beta_{\vartheta})$ determines whether the current option should terminate or not.
Option Mask and Finite-State Structure
@@ III-D2 Option Decoder $\pi_{\vartheta}$
[…]
Figure 3: Comparison of travel distances on simulated maps. For each planner, each bar encodes three statistics: the top, middle, and bottom markers correspond to the makespan, the average travel distance, and the minimum distance within the team, respectively. Results are shown for teams of 33, 44, 55, and 1010 agents.
III-E Multi-Agent Critic Networks
III-E1 Multi-Agent Centralized Critic
For agent $i$, let $\mathbf{h}_{i}$ and $\tilde{\mathbf{h}}_{i}$ denote its local and enhanced node features, and let $\tilde{\mathbf{h}}_{i,j}$ be the embedding of robot $j$ in the team. Its local state–action embedding is $\mathbf{u}_{i,j}=f_{\mathrm{emb}}([\tilde{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i,j}])$.
Following MAAC [7], the critic computes a query $\mathbf{q}_{i}=W_{q}\mathbf{h}_{i}$ and attends to keys and values $\mathbf{k}_{j}=W_{k}\mathbf{u}_{j}$, $\mathbf{v}_{j}=W_{v}\mathbf{u}_{j}$ via
Finally, the critic forms
$\mathbf{u}_{i,j}=[\mathbf{h}_{i},\tilde{\mathbf{h}}_{i},\mathbf{c}_{i},\tilde{\mathbf{h}}_{i,j}],\qquad Q_{i}=f_{Q}([\mathbf{u}_{i,j}]_{j\in\mathcal{N}}),$
where $f_{Q}$ is the embedding layer to output the centralized action-value estimation $Q_{i}$ for agent $i$ in the robot team $\mathcal{N}$, which integrates local features, multi-agent interactions, and action-specific neighborhood structure.
III-E2 Option-Regularized Critic Learning
Our critic network adopts the same graph-transformer backbone as the policy network, but differs in its use of a privileged graph encoder. During training, this encoder has access to privileged ground-truth information (i.e., the ground-truth map), enabling the critic to provide low-variance estimations on long-term returns to the decentralized policies. Unlike the actor’s dual encoders (one for the combined graph and one for the current graph), the critic employs a single unified encoder that captures the complete relational context of all agents and their robot beliefs. Besides the state representation input, the critic also takes the option feature selected at the previous step as input, which benefits both option and waypoint optimization during critic learning [7]. The value estimate is explicitly conditioned on the agent’s previous option $z_{t-1}$, representing the high-level mode active at the current step. Specifically, the option embedding is added to the current node feature:
$\mathbf{h}_{t}^{i}=\mathrm{Gather}(H_{t},\text{index}=i)+e(z_{t-1}^{i}),$ (2)
Termination Head Loss and Policy Loss
Our critic network enables termination learning by contrasting/advantaging the expected return of continuing the current option with that of terminating it. For agent $i$, given the per-agent Q-values $Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},z_{t})$ and the counterfactual value if the option terminates $Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},\bar{z_{t}})$, we compute the advantage of termination as
$A_{\text{term}}^{i}=Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},\bar{z_{t}})-Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},z_{t}),$ (3)
which quantifies whether switching options yields a higher expected value than continuing.
In the policy network’s option decoder, the termination head outputs $\mathbf{1}_{\text{term}}/\mathbf{1}_{\text{cont}}\sim\mathrm{Bern}(\beta_{\vartheta})$ through a scalar logit, from which we define the Bernoulli probabilities of termination and continuation:
$p_{\text{term}}^{i}\sim\log\beta_{\vartheta}(\cdot\mid\mathbf{s}_{t},z_{t-1}),\quad p_{\text{cont}}^{i}\sim\log\beta_{\vartheta}(\cdot\mid\mathbf{s}_{t},\bar{z}_{t-1}).$ (4)
Given the binary termination signal $\delta_{t}^{i}\in\{0,1\}$ from the termination head, the likelihood of the observed decision is
$p_{\vartheta}^{i}=\mathbf{1}_{\text{term}}p_{\text{term}}^{i}+\mathbf{1}_{\text{cont}}p_{\text{cont}}^{i}.$ (5)
Overall, we optimize the termination head by maximizing the expected return differential, i.e., agents are encouraged to terminate when doing so leads to higher option-state values. The termination head loss is thus defined as
$\mathcal{L}_{\vartheta}=-\,\mathbb{E}\!\left[\log p_{\vartheta}^{i}\,A_{\text{term}}^{i}\right],$ (6)
On the other hand, following [5], the policy loss is further improved by our option-regularized policy network:
Finally, the overall loss is $\mathcal{L}_{\text{policy}}=\mathcal{L}_{\pi}+\lambda_{\vartheta}\,\mathcal{L}_{\vartheta},$ where $\lambda_{\vartheta}$ balances termination learning. Following [5], the critic is optimized via a temporal-difference (TD) regression objective.
@@ III-E2 Option-Regularized Critic Learning
[…]
@@
[…]
Figure 4: ROS Experiments. ORION yields more efficient decentralized coordination, as seen from each agent’s local maps and the shared map, compared to ORION w/o option.
Figure 5: Runtime performance in Gazebo simulation. ORION maintains real-time performance on graph updates (green) and network inference (blue) throughout execution, while the prior utility curve (orange) reflects how ORION incrementally verifies and corrects uncertain regions in the prior map during online navigation.
@@ IV-A Comparisons in Simulated Maps
Inspired by recent navigation and exploration benchmarks [16, 23], we introduce a new benchmark comprising hundreds of large, maze-like, and challenging simulated maps, in which we evaluate ORION against representative baselines under partially known environments. For the learning-based baseline, we consider MAContext [16], a multi-agent extension of a DRL navigation planner designed for partially known environments. For classical MAPF solvers, we include 1) LNS2 [13], a highly efficient and scalable planner that iteratively repairs collision-containing path sets by replanning for a subset of colliding agents, and 2) EECBS [14], a bounded-suboptimal MAPF planner that combines Explicit Estimation Search at the high level with focal search at the low level. Since both LNS2 and EECBS operate on static maps, we deploy them in our partially known setting by triggering replanning whenever new changes are observed by any agent in the environment. To further analyze the contribution of our design, we additionally evaluate two ablation variants: 1) ORION w/o option, removing our option-critic structure, and 2) ORION w/o dual-stage, disabling our pre-/post-arrival stages. These variants help isolate the role of our proposed option-critic’s hierarchical decision-making and dual-stage cooperation strategy.
As shown in Table I, ORION achieves the best overall performance across all team sizes, reducing makespan by up to $20\text{\,}\%$ compared to state-of-the-art planners. Traditional baselines (LNS2 and EECBS) maintain strong performance as the team size grows; however, ORION consistently outperforms them, exhibiting a $12.5\text{\,}$$13.4\text{\,}\%$ advantage at 3 agents and still retaining $6.5\text{\,}$$6.9\text{\,}\%$ improvement at 10 agents. In contrast, the learning-based baseline MAContext shows clear degradation with larger teams: its makespan worsens from $478.5\text{\,}\mathrm{m}$ to $500.8\text{\,}\mathrm{m}$ when scaling from 5 to 10 agents (around a $4.6\text{\,}\%$ increase). Meanwhile, ORION maintains stable makespan and decision steps as the team size increases, highlighting its superior scalability and generalizability in cooperative multi-agent navigation. Although our task objective focuses on minimizing the team’s makespan, ORION also demonstrates great improvements in average and minimum travel distance (see Fig. 3), further indicating its overall planning efficiency. Beyond these quantitative results, we observe qualitative behaviors that help explain ORION’s superior performance. First, agents that reach their targets early, or whose routes are less sensitive to map discrepancies, often make small, deliberate detours to reduce key uncertainties for the team, trading a minor loss in individual optimality for a larger team-level gain. Second, this cooperation is not evenly distributed across the team; instead, ORION tends to concentrate help on the bottleneck agents whose prior-map plans are most disrupted by ground-truth changes, effectively alleviating the team’s slowest members. Finally, the option-regularized policy enables timing-aware cooperation: rather than exploring blindly from the start or having every agent help at all times, agents intervene when assistance is most valuable, typically after reaching their own targets or when the marginal cost of detouring is low.
@@
Figure 6: Real-world experiments. Two start–target settings are tested. For each case, the left column shows the combined map and trajectories, while the right column shows each agent’s current belief and executed path.
@@ IV-B Validation in Large-Scale Gazebo Simulations
We further evaluate our approach in a $70\text{\,}\mathrm{m}$$\times$$60\text{\,}\mathrm{m}$ Gazebo warehouse environment, where shelf rearrangements cause partial discrepancies between the prior map and the ground truth, thereby simulating realistic online navigation under map changes. ORION and MAContext are deployed on ground vehicle teams (maximum speed $2\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$) equipped with Velodyne VLP16 LiDARs. As shown in Table II, we compare ORION, ORION w/o option-critic, and MAContext with teams of 3, 4, and 5 robots. ORION consistently achieves the best performance in terms of maximum and average distance. While ORION exhibits slightly larger $D_{min}$ for the largest team, this reflects deliberate post-arrival exploration: agents that reach their targets early, or whose paths are less affected by map changes, take short detours to reduce uncertainty for the most affected (slowest) agents. This strategic trade-off slightly increases individual distance but yields a substantially larger team-level gain, emerging naturally from the option-critic formulation and the team-level objective rather than from hard-coded heuristics. Fig. 4 illustrates a representative cooperative behavior enabled by ORION: agent 1 reaches its target first (red trajectory) and then moves toward the vicinity of target 2 to share informative observations with agent 2. Upon completing this assistance, agent 1 returns to its own target instead of performing unnecessary exploration, highlighting adaptive cooperation enabled by the option-critic framework and the proposed dual-stage strategy.
As shown in Fig. 5, ORION maintains real-time performance throughout the experiments: graph update remains below $0.5\text{\,}\mathrm{s}$, and network inference requires less than $0.2\text{\,}\mathrm{s}$ (typically under $0.1\text{\,}\mathrm{s}$), despite the entire system being implemented in Python. We also report the evolution of prior utility, defined as the number of sampled nodes along the boundary between confirmed free space and uncertain obstacles. The prior utility stabilizes after approximately $40\text{\,}\mathrm{s}$, indicating that the team has obtained a sufficiently confident map and no longer needs extra exploration.

### v3

@@ III-C1 Graph-based Encoders
III-C2 Option Decoder $\pi_{\vartheta}$
During multi-agent navigation, an agent may either navigate toward its designated target or assist others by collecting more information that benefits the team, each demanding distinct and sometimes conflicting reasoning. In such cases, adaptive switching between self-directed and cooperative behaviors becomes crucial. To balance individual efficiency and collective performance for improved overall task success, we introduce the option-critic structure [2], where an option denotes a high-level behavioral mode characterized by a termination head function and an intra-option policy, while the critic provides a value-based estimation guiding high-level option/low-level action optimization. Specifically, we augment the policy network with an option decoder, which determines whether to continue the current high-level behavior or terminate and switch to another. Given the current state representation feature $\mathbf{s}_{t}$ and the active option $z_{t-1}$ at the last step, we define a termination head function $\beta_{\vartheta}(\mathbf{s}_{t},z_{t-1})$ that estimates the likelihood of terminating the current option at step $t$. Specifically, we first obtain a joint feature representation by combining the current node feature and the embedding of the previous option $h_{t}=\mathbf{s}_{t}+e(z_{t-1}),$ where $e(z_{t-1})\in\mathbb{R}^{d}$ denotes the learnable option embedding. We parameterize the termination function as $\beta_{\vartheta}(\mathbf{s}_{t},z_{t-1})=\sigma\!\big(\mathrm{MLP}_{\vartheta}(h_{t})\big),$ where $\mathrm{MLP}_{\vartheta}(\cdot)$ is a two-layer feed-forward network with ReLU activation, and $\sigma(\cdot)$ denotes the logistic sigmoid. Finally, a Bernoulli sampling $\mathbf{1}_{\text{term}}/\mathbf{1}_{\text{cont}}\sim\mathrm{Bern}(\beta_{\vartheta})$ determines whether the current option should terminate or not.
@@ III-C2 Option Decoder $\pi_{\vartheta}$
III-C3 Multi-Agent Critic Networks
For agent $i$, local feature $\mathbf{h}_{i}$ is extracted from the critic encoder to capture individual observations without inter-agent interaction. Enhanced feature $\tilde{\mathbf{h}}_{i}$ further aggregates team-wide context via cross-attention layers. To model coordination, $\tilde{\mathbf{h}}_{i,j}$ denotes the relational embedding between agent $i$ and $j$, capturing their relative states under a shared team belief. These features are concatenated to form the local state–action embedding $\mathbf{u}_{i,j}=f_{\mathrm{emb}}([\tilde{\mathbf{h}}_{i},\tilde{\mathbf{h}}_{i,j}])$, which jointly integrates the agent’s local state, inter-agent relationships, and the global context of the environment. Following MAAC [10], the critic computes a query $\mathbf{q}_{i}=W_{q}\mathbf{h}_{i}$ and attends to keys and values $\mathbf{k}_{j}=W_{k}\mathbf{u}_{i,j}$, $\mathbf{v}_{j}=W_{v}\mathbf{u}_{i,j}$ via
Finally, the critic forms $Q_{i}=f_{Q}([\mathbf{u}_{i,j}]_{j\in\mathcal{N}}),$ where $f_{Q}$ is the embedding layer to output the centralized action-value estimation $Q_{i}$ for agent $i$ in the robot team $\mathcal{N}$, which integrates local features, multi-agent interactions, and action-specific neighborhood structure. Our critic network adopts the same graph-transformer backbone as the policy network, but differs in its use of a privileged graph encoder. During training, this encoder has access to privileged ground-truth information (i.e., the ground-truth map), enabling the critic to provide low-variance estimations on long-term returns to the decentralized policies. Unlike the actor’s dual encoders (one for the combined graph and one for the current graph), the critic employs a single unified encoder that captures the complete relational context of all agents and their robot beliefs. Besides the state representation input, the critic also takes the option feature selected at the previous step as input, which benefits both option and waypoint optimization during critic learning [10]. The value estimate is explicitly conditioned on the agent’s previous option $z_{t-1}$, representing the high-level mode active at the current step. Specifically, the option embedding is added to the current node feature:
$\mathbf{h}_{t}^{i}=\mathrm{Gather}(H_{t},\text{index}=i)+e(z_{t-1}^{i}),$ (2)
Our critic network enables termination learning by contrasting/advantaging the expected return of continuing the current option with that of terminating it. For agent $i$, given the per-agent Q-values $Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},z_{t})$ and the counterfactual value if the option terminates $Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},\bar{z_{t}})$, we compute the advantage of termination as $A_{\text{term}}^{i}=Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},\bar{z_{t}})-Q_{\vartheta}^{i}(\mathbf{s}_{t},a_{t},z_{t}),$ which quantifies whether switching options yields a higher expected value than continuing. In the policy network’s option decoder, the termination head outputs $\mathbf{1}_{\text{term}}/\mathbf{1}_{\text{cont}}\sim\mathrm{Bern}(\beta_{\vartheta})$ through a scalar logit, from which we define the Bernoulli probabilities of termination and continuation: $p_{\text{term}}^{i}\sim\log\beta_{\vartheta}(\cdot\mid\mathbf{s}_{t},z_{t-1})$ and $p_{\text{cont}}^{i}\sim\log\beta_{\vartheta}(\cdot\mid\mathbf{s}_{t},\bar{z}_{t-1})$. Given the binary termination signal $\delta_{t}^{i}\in\{0,1\}$ from the termination head, the likelihood of the observed decision is $p_{\vartheta}^{i}=\mathbf{1}_{\text{term}}p_{\text{term}}^{i}+\mathbf{1}_{\text{cont}}p_{\text{cont}}^{i}$.
Overall, we optimize the termination head by maximizing the expected return differential, i.e., agents are encouraged to terminate when doing so leads to higher option-state values. The termination head loss is thus defined as $\mathcal{L}_{\vartheta}=-\,\mathbb{E}\!\left[\log p_{\vartheta}^{i}\,A_{\text{term}}^{i}\right]$. On the other hand, following [8], the policy loss is further improved by our option-regularized policy network:
Finally, the overall loss is $\mathcal{L}_{\text{policy}}=\mathcal{L}_{\pi}+\lambda_{\vartheta}\,\mathcal{L}_{\vartheta},$ where $\lambda_{\vartheta}$ balances termination learning. Following [8], the critic is optimized via a temporal-difference regression objective.
@@ III-C3 Multi-Agent Critic Networks
ORION 455.83 ($\pm$108.51) 55.36 441.99 ($\pm$107.2) 53.86 439.24 ($\pm$111.08) 53.24 435.76 ($\pm$99.28) 53.26
@@
Max Dist $\downarrow$ Avg Dist $\downarrow$ Min Dist $\downarrow$ Max Dist $\downarrow$ Avg Dist $\downarrow$ Min Dist $\downarrow$ Max Dist $\downarrow$ Avg Dist $\downarrow$ Min Dist $\downarrow$
@@
ORION 75.99 64.2 55.34 72.68 63.5 54.14 95.64 75.06 64.58
@@ IV-A Comparisons in Simulated Maps
We expand the navigation dataset released in [19] to train ORION with a sensor range of $20\text{\,}\mathrm{m}$. For graph construction, we uniformly sample $600\text{\,}$ points across the environment and treat points in known free space as candidate viewpoints. Each viewpoint is connected to its $20\text{\,}$ nearest neighbors while retaining only collision-free edges. ORION is trained on a workstation with one i9-10980XE CPU and two NVIDIA GeForce RTX 3090 GPUs. Training typically converges after $60\,000\text{\,}$ episodes and takes around one day. Following recent navigation and exploration benchmarks [19, 26], we further introduce a benchmark of $500\text{\,}$ large and challenging maze-like simulated maps to evaluate ORION against representative baselines. For the learning-based baseline, we consider MAContext [19], a multi-agent extension of a DRL navigation planner designed for partially known environments. For classical MAPF solvers, we include 1) LNS2 [16], a highly efficient and scalable planner that iteratively repairs collision-containing path sets by replanning for a subset of colliding agents, and 2) EECBS [17], a bounded-suboptimal MAPF planner that combines high-level Explicit Estimation Search with low-level focal search. To deploy LNS2 and EECBS in our partially known setting, we align their execution with ORION by using the same graph-construction pipeline: At each decision timestep, they replan on the constructed graph and select the first waypoint of the planned path as the next waypoint, with the same replanning frequency and collision definitions as ORION. To further analyze our model, we additionally evaluate four ablation variants: 1) ORION w/o option, removing our option-critic structure, 2) ORION w/o current, current map being unavailable, 3) ORION w/o combined, excluding the combined map, and 4) ORION w/o dual-stage, disabling our pre-/post-arrival stages. These variants help isolate the role of our proposed option-critic’s hierarchical decision-making, dual map representations, and dual-stage cooperation strategy.
As shown in Table I, ORION achieves the best overall performance across all team sizes, reducing makespan by up to $20\text{\,}\%$ over state-of-the-art planners. Traditional baselines (LNS2 and EECBS) maintain competitive as the team grows; however, ORION consistently outperforms them, exhibiting a $12.5\text{\,}$$13.4\text{\,}\%$ advantage at 3 agents and still retaining $6.5\text{\,}$$6.9\text{\,}\%$ improvement at 10 agents. In contrast, the learning-based baseline MAContext degrades with scale, with makespan increasing from $478.5\text{\,}\mathrm{m}$ to $500.8\text{\,}\mathrm{m}$ as the number of agents increases from 5 to 10. ORION, however, maintains stable makespan and decision steps, highlighting stronger scalability and generalization. Although our task objective focuses on minimizing the team’s makespan, ORION also improves the average and minimum travel distance (Fig. 3), further indicating better overall planning efficiency. Qualitatively, ORION exhibits targeted, timing-aware cooperation: agents with low-cost detours or completed goals often help resolve key uncertainties for bottleneck teammates rather than assisting uniformly across the team. Two ablations further indicate that both maps are necessary. ORION w/o combined, i.e., without prior information, causes a clear makespan drop ($7.5\text{\,}$$9.7\text{\,}\%$ at 3–5 agents), since agents spend more effort exploring before exploiting prior observations. Removing the current map (ORION w/o current) also degrades performance across all team sizes, showing that up-to-date observations are equally important for correcting prior-map errors. Together, these results confirm that the combined map provides useful environmental priors, while the current map supplies accurate online updates.
@@
Figure 3: Comparison of travel distances on simulated maps. For each planner, each bar shows the makespan, average travel distance, and minimum team distance as the top, middle, and bottom markers, respectively. Results are reported for teams of 33, 44, 55, and 1010 agents.
Figure 4: Gazebo Experiments. ORION yields more efficient decentralized coordination than ORION w/o option, as shown by the local maps of the four agents.
Figure 5: Runtime performance in Gazebo simulation. ORION maintains real-time performance on graph updates and network inference throughout execution, while the prior frontier curve reflects how ORION incrementally verifies and corrects uncertain regions in the prior map during online navigation.
Figure 6: Real-world experiments under two start–target settings. The left column shows the combined map and trajectories, while the right column shows each agent’s current belief and executed path.
@@ IV-B Validation in Large-Scale Gazebo Simulations
We further evaluate our approach in a $70\text{\,}\mathrm{m}$$\times$$60\text{\,}\mathrm{m}$ Gazebo warehouse environment, where shelf rearrangements cause partial discrepancies between the prior map and the ground truth, thereby simulating realistic online navigation under map changes. ORION and MAContext are deployed on ground vehicle teams (maximum speed $2\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$) equipped with Velodyne VLP16 LiDARs. As shown in Table II, we compare ORION, ORION w/o option-critic, and MAContext with teams of 3, 4, and 5 robots. ORION consistently achieves the best performance in terms of maximum and average distance. We also observe that, with the option module enabled, ORION brings a performance improvement of $12\text{\,}$$32\text{\,}\%$ in maximum distance and $19\text{\,}$$34\text{\,}\%$ in average distance, which further highlights the importance of the option module. While ORION exhibits a slightly larger minimum distance for the largest team, this reflects deliberate post-arrival exploration. Agents that reach their targets early, or are less affected by map changes, take short detours to reduce uncertainty for slower teammates. This trade-off slightly increases individual distance but provides larger team-level gains. It emerges naturally from the option-critic formulation and team-level objective rather than hard-coded heuristics.
Fig. 4 illustrates a representative cooperative behavior enabled by ORION: agent 1 reaches its target first (red trajectory) and then moves toward the vicinity of target 2 to share informative observations with agent 2. Upon completing this assistance, agent 1 returns to its own target instead of performing unnecessary exploration, highlighting adaptive cooperation enabled by the option-critic framework and the proposed dual-stage strategy. As shown in Fig. 5, ORION maintains real-time performance throughout the experiments: graph update remains below $0.5\text{\,}\mathrm{s}$, and network inference requires less than $0.2\text{\,}\mathrm{s}$ (typically under $0.1\text{\,}\mathrm{s}$), despite the entire system being implemented in Python. We also report the evolution of prior utility, defined as the number of sampled nodes along the boundary between confirmed free space and uncertain obstacles. The prior utility stabilizes after approximately $40\text{\,}\mathrm{s}$, indicating that the team has obtained a sufficiently confident map and no longer needs extra exploration.